OpenStudy (anonymous):

what is the midpoint of ab when a= (-2,-3) and b= (7,-12)

OpenStudy (campbell_st):

x = (-2 + 7)/2 y=(-3 + -12)/2 basically you average the x and y values to find the midpoint

OpenStudy (gorv):

mid point means which is at equal distance from both point

OpenStudy (gorv):

for x co ordinate add x value and divide by 2

OpenStudy (gorv):

similar for y

OpenStudy (astrophysics):

\[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\] this is the formula, basically what campbell said.